# Thermodynamics and Gravitational Signatures of Rotating Black Holes in the Generalized Extended Uncertainty Principle

### v1

@@ VI Conclusion
Regarding the dynamical stability, we utilized the Teukolsky formalism to study gravitational perturbations. A key theoretical finding is that the rotating GEUP black hole preserves the isospectrality between axial and polar modes. This result highlights the nature of the GEUP deformation as ”parametric” rather than ”structural.” Unlike modified gravity theories such as Einstein-Dilaton-Gauss-Bonnet (EDGB) or Dynamical Chern-Simons (dCS) gravity, where the coupling to scalar fields or parity-violating terms breaks the degeneracy between the even and odd parity sectors, the GEUP framework effectively renormalizes the background mass scale while preserving the Petrov Type D character of the spacetime. Consequently, the quantum corrections manifest as coherent shifts in the resonant frequencies, rather than a splitting of the isospectral multiplet.
Finally, we confronted the model with recent observational data from LIGO/Virgo and the Event Horizon Telescope (EHT). Our analysis highlights a fundamental complementarity in black hole phenomenology:
  • Gravitational Waves: We showed that stellar-mass merger remnants (like GW150914) are largely insensitive to EUP corrections but provide the tightest available astrophysical bounds on the GUP parameter $\beta$.
  • Black Hole Shadows: Conversely, we derived that the shadow radius scales linearly with the effective mass, $R_{sh}\approx 3\sqrt{3}\mathcal{M}$. Comparing constraints from M87* and Sgr A*, we established that M87* is approximately $10^{6}$ times more sensitive to large-scale modifications than Sgr A*. Specifically, the ratio of sensitivity scales as:
    $\frac{\delta_{M87}}{\delta_{SgrA}}\approx\left(\frac{M_{M87}}{M_{SgrA}}\right)^{2}\approx 2.6\times 10^{6}.$ (66)
    Consequently, the consistency of the M87* shadow with GR places the definitive upper bound on the EUP parameter $\alpha$, confirming that $\mathcal{M}\approx M$ in the astrophysical regime.
In summary, the Rotating GEUP metric provides a consistent and testable phenomenological model of quantum gravity. It resolves the ”tension” between small-scale and large-scale modifications by mapping them onto distinct observational windows—using shadows to constrain the infrared sector and gravitational waves to constrain the ultraviolet sector. Future work will extend this analysis to massive scalar field perturbations and the detailed structure of the photon ring autocorrelations.

### v2

@@ VI Conclusion
Regarding the dynamical stability, we utilized the Teukolsky formalism to study gravitational perturbations. A key theoretical finding is that the rotating GEUP black hole preserves the isospectrality between axial and polar modes. This result highlights the nature of the GEUP deformation as “parametric” rather than “structural.” Unlike modified gravity theories such as Einstein-Dilaton-Gauss-Bonnet (EDGB) or Dynamical Chern-Simons (dCS) gravity, where the coupling to scalar fields or parity-violating terms breaks the degeneracy between the even and odd parity sectors, the GEUP framework effectively renormalizes the background mass scale while preserving the Petrov Type D character of the spacetime. Consequently, the quantum corrections manifest as coherent shifts in the resonant frequencies, rather than a splitting of the isospectral multiplet.
Finally, we confronted the model with recent observational data from LIGO/Virgo and the Event Horizon Telescope (EHT). Our analysis highlights a fundamental duality in the phenomenology:
  • Thermodynamic Distinguishability: The most significant deviation occurs in the evaporation history. We established that the “cryogenic” phase ($T_{H}\propto M^{-3}$) creates a distinct dynamical signature: a GEUP black hole will survive significantly longer than a standard Hawking black hole. This suggests that the true test of the theory lies in cosmological timescales rather than static snapshots.
  • Geometric Indistinguishability: Conversely, we demonstrated that static geometric observables, such as the shadow radius and inspiral chirp mass, are degenerate. Since gravitational measurements probe the effective ADM mass $\mathcal{M}$ rather than the bare mass $M$, a GEUP black hole mimics a standard Kerr black hole of a different mass. Consequently, EHT observations of M87* and Sgr A* cannot constrain the EUP parameter $\alpha$ without an independent determination of the bare corpuscular mass. Instead, the agreement with EHT data confirms the phenomenological consistency of the theory, proving that large-scale quantum corrections do not introduce exotic visual artifacts.
In summary, the Rotating GEUP metric provides a consistent model of quantum gravity that resolves the observability problem via time-evolution. While the static geometry is protected by mass renormalization, the thermodynamic trajectory is fundamentally altered. Future work will extend this analysis to massive scalar field perturbations and the detailed structure of the photon ring autocorrelations to identify potential second-order deviations.